As you stand by the side of the road, a car approaches you at a constant speed, sounding its horn, and you hear
a frequency of 80.0 Hz. After the car goes by, you hear a frequency of 60.0 Hz. What is the speed of the car?
Assume that he speed of sound in air is 343 m/s.
Answer #1
80 = f/ (1- u/v)
60 = f/ (1+ u/v)

equating f from above 2 equations we get

80* (1- u/v) = 60* (1+ u/v)

and given that v = 343

so we find that u = 343/7 = 49m/s
Answer #2

80 = f/ (1- u/v)
60 = f/ (1+ u/v)

equating f from above 2 equations we get

80* (1- u/v) = 60* (1+ u/v)

and given that v = 343

so we find that u = 343/7 = 49m/s